# An Order of Magnitude Time Complexity Reduction for Gaussian Graphical Model Posterior Sampling Using a Reverse Telescoping Block Decomposition

### v1

@@ Abstract
We consider the problem of fully Bayesian posterior estimation and uncertainty quantification in undirected Gaussian graphical models via Markov chain Monte Carlo (MCMC) under recently-developed element-wise graphical priors, such as the graphical horseshoe. Unlike the conjugate Wishart family, these priors are non-conjugate; but have the advantage that they naturally allow one to encode a prior belief of sparsity in the off-diagonal elements of the precision matrix, without imposing a structure on the entire matrix. Unfortunately, for a graph with $p$ nodes and with $n$ samples, the state-of-the-art MCMC approaches for the element-wise priors achieve a per iteration complexity of $\mathcal{O}(p^{4}),$ which is prohibitive when $p\gg n$. In this regime, we develop a suitably reparameterized MCMC with per iteration complexity of $\mathcal{O}(p^{3})$, providing a one-order of magnitude improvement, and consequently bringing the computational cost at par with the conjugate Wishart family, which is also $\mathcal{O}(p^{3})$ due to a use of the classical Bartlett decomposition, but this decomposition does not apply outside the Wishart family. Importantly, the proposed benefit is obtained solely due to our reparameterization in an MCMC scheme targeting the true posterior, that reverses the recently developed telescoping block decomposition of Bhadra et al., (2024), in a suitable sense. There is no variational or any other approximate Bayesian computation scheme considered in this paper that compromises targeting the true posterior. Simulations and the analysis of a breast cancer data set confirm both the correctness and better algorithmic scaling of the proposed reverse telescoping sampler.
@@ 1 Introduction
We denote by $p$ the number of variables at hand and by $n$ the sample size. Our specific focus in this article is on Bayesian estimation of $\boldsymbol{\Theta}=\{\theta_{jk}\}$ in the regime $p\gg n$. To begin, we note that this setting is computationally challenging even for routine priors, such as the conjugate Wishart family. Assume one observes $\mathbf{y}\sim\mathcal{N}(0,\mathbf{I}_{n}\otimes\boldsymbol{\Theta}^{-1})$ and the prior on $\boldsymbol{\Theta}$ is $\mathcal{W}_{p}(b,\boldsymbol{\Phi})$ for some b>(p1)b>(p-1) and $\boldsymbol{\Phi}\in\mathcal{M}_{p}^{+}$, where $\mathcal{M}_{p}^{+}$ denotes the space of $p\times p$ positive definite matrices, and $\mathcal{W}_{p}(b,\boldsymbol{\Phi})$ denotes a Wishart random variable with shape parameter $b$ and scale matrix $\boldsymbol{\Phi}$ (Gupta and Nagar,, 2000, Chapter 3). Then, by routine calculations, the posterior of $\boldsymbol{\Theta}$ is also Wishart, given as: $\boldsymbol{\Theta}\mid\mathbf{y}\sim\mathcal{W}_{p}(b+n,\;\boldsymbol{\Phi}+\mathbf{y}^{T}\mathbf{y})$. In this setting, the standard algorithm for posterior sampling is Bartlett’s decomposition of Wishart (Bartlett,, 1934), which requires a Cholesky decomposition of a $p\times p$ matrix, and thus has a computational cost of $\mathcal{O}(p^{3})$. Unfortunately, the situation is worse for non-conjugate priors. A crucial bottleneck in the posterior sampling of $(\boldsymbol{\Theta}\mid\mathbf{y})$ under any prior is that one must ensure the posterior samples also belong to $\mathcal{M}_{p}^{+}$. This restriction is automatically satisfied in the conjugate Wishart family, which puts mass only on the space of positive definite matrices, but must be enforced explicitly under all other non-conjugate priors. Therefore, unlike in regression problems with real-valued coefficients, it is far more challenging to put element-wise priors on $\theta_{jk}$ and then to design a valid posterior sampler respecting the constraint $\boldsymbol{\Theta}\in\mathcal{M}_{p}^{+}$. A naïve accept–reject sampler is not likely to be effective in enforcing this constraint at a large enough $p$, because the space $\mathcal{M}_{p}^{+}$ constitutes a very small subset of $\mathbb{R}^{p\times p}$. A Cholesky-based modeling approach, where one writes $\boldsymbol{\Theta}=\mathbf{L}\mathbf{L}^{T}$, is a possibility for maintaining positive definiteness while allowing for unconstrained modeling of $\mathbf{L}$ (Pourahmadi,, 1999; Daniels and Pourahmadi,, 2002), but again, a prior on $\mathbf{L}$ in general induces a complicated prior structure on $\boldsymbol{\Theta}$, and vice versa, apart from special cases such as the Wishart or G-Wishart (Roverato,, 2002; Uhler et al.,, 2018). Thus, if one desires a sparsity-inducing prior directly on $\theta_{jk}$ to learn a sparse graphical structure, it is unclear what induced prior on $\mathbf{L}$ would achieve this goal, apart from very specific cases.
@@
An ingenious approach that indeed allows one to put element-wise non-conjugate priors on $\theta_{jk}$ and still perform Gibbs sampling while respecting the positive definite constraint, was developed by Wang, (2012) in the context of the Bayesian graphical lasso (BGL) prior. We describe this in detail in Section 2. However the parameteization of Wang, (2012) is not specific to the BGL, and similar sampling approaches have been used later for other priors with better statistical properties, such as the graphical horseshoe or GHS (Li et al.,, 2019), the graphical horseshoe-like or GHSL (Sagar et al., 2024a, ), and the graphical R2-D2 (Gan et al.,, 2022), among others. In particular, Sagar et al., 2024a studied the posterior concentration properties of the element-wise graphical priors and showed that the GHS and GHSL priors enjoy the optimal rates of posterior contraction established by Banerjee and Ghosal, (2014). The penalized maximum a posteriori (MAP) estimate under these element-wise priors have also been studied. For the BGL prior, the MAP corresponds to the frequentist graphical lasso (Friedman et al.,, 2008), whereas a novel local linear approximation scheme for MAP estimation under generic element-wise priors was developed by Sagar et al., 2024b . While MAP estimation is faster, it does not provide full posterior uncertainty quantification, and we do not discuss this further in the current work. Nevertheless, the decomposition of Wang, (2012) is also fundamental to the MAP approaches (Wang,, 2014; Sagar et al., 2024b, ).
@@
Despite its utility, Wang’s parametrization has a fundamental bottleneck, which is a $\mathcal{O}(p^{4})$ complexity per iteration to generate a sample of $(\boldsymbol{\Theta}\mid\mathbf{y})$, which is one order-of magnitude worse than the $\mathcal{O}(p^{3})$ complexity for the conjugate Wishart model, and is prohibitive when $p$ is large. We describe in detail in Section 2 why this complexity is inherent to Wang’s approach, and there is no way to bypass this, even if $n$ is much smaller than $p$. Intuitively, this happens because Wang, (2012)’s estimation procedure only sees the $p\times p$ scatter matrix $\mathbf{S}=\mathbf{y}^{T}\mathbf{y}$, defined as $n$-times the sample covariance matrix, and not the $n\times p$ data matrix $\mathbf{y}$ itself. Thus, even if $n$ is much smaller than $p$, there is no way for Wang, (2012) to take advantage of this.
@@
In this work, we demonstrate that a valid Gibbs sampler targeting the exact posterior is possible with a worst case per iteration complexity of $\mathcal{O}(p^{3})$ when $p\gg n$, providing a one order of magnitude reduction in computation cost, and bringing the cost down at par with the conjugate Wishart family, even for non-conjugate models. The key to our work is to reverse, in a suitable sense, the telescoping block decomposition of the precision matrix recently developed by Bhadra et al., (2024) for estimating the marginal likelihood in GGMs under element-wise priors. However, the possibility of using this decomposition for sampling was not developed by Bhadra et al., (2024), who end up using a $\mathcal{O}(p^{5})$ algorithm. As we demonstrate in the current work, this decomposition is also particularly amenable to fast sampling under Gaussian scale mixture priors, developed in the context of linear regression problems by Bhattacharya et al., (2016). Using this, we demonstrate that our sampler has a per iteration complexity of $\mathcal{O}(\max{(n^{2}p^{2},p^{3})})$, which is $\mathcal{O}(p^{3})$ in the $p\gg n$ regime, i.e., if $n=\mathdutchcal{o}(p^{1/2})$. We take particular care to distinguish our methodology with existing approaches that also use a regression setting, such as the Cholesky factor based methods (e.g. Pourahmadi,, 1999) or pseudolikelihood methods (e.g. Meinshausen and Bühlmann,, 2006). We perform extensive simulations under the GHS and GHSL priors to verify both the correctness of our algorithm, and its improved scaling over $p$.
@@ 2.3 Model Posterior and the Cyclical Gibbs Sampler of Wang, (2012)
    For the GHS prior of Li et al., (2019), $\lambda_{jk}\sim C^{+}(0,1)$, a priori, where $C^{+}$ denotes the half Cauchy density. However, introducing one further level of augmentation, one can show: $\lambda_{jk}\sim C^{+}(0,1)$ is equivalent to $\lambda_{jk}^{2}\mid a\sim\mathrm{IG}(1/2,1/a)$ and $a\sim\mathrm{IG}(1/2,1)$, where IG denotes an inverse gamma random variable in shape–scale parameterization. This leads to Gibbs sampling from the posterior as in Makalic and Schmidt, (2016), with $(\lambda^{2}_{jk}\mid\mathrm{rest})$ and $(a\mid\mathrm{rest})$ both following inverse gamma distributions.
@@
Similar strategies are available for posterior sampling under the GHSL prior (Sagar et al., 2024a, ), the graphical R2-D2 prior (Gan et al.,, 2022) and several others. However, it is important to note that the positive definite constraint on $\boldsymbol{\Theta}$ does not affect the sampling of these parameters in a Gibbs step, which is conditional on the current $\boldsymbol{\Theta}$. Finally, for the global $\tau$, a standard half-Cauchy prior is typical (Li et al.,, 2019).
@@ 3.1 The Reverse Telescoping Block Decomposition
Remark 2 (The notations $\boldsymbol{\Theta}$ vs. $\boldsymbol{\Omega}$).
We use the notation $\boldsymbol{\Theta}$ and $\boldsymbol{\tilde{\Theta}}$ to denote the matrices, including all their submatrices, that remain unchanged throughout the telescoping and reverse telescoping adjustments and all sampling steps; whereas the notations $\boldsymbol{\Omega}$ and $\boldsymbol{\tilde{\Omega}}$ denote placeholder matrices whose upper-left $(j\times j)$ blocks are successively overwritten as the algorithms proceed over $j=(p-1),\ldots,1$. We clarify this distinction using keywords “Set” vs. “Update” in our algorithm description. This allows us to write $\tilde{\boldsymbol{\theta}}_{j}=(\boldsymbol{\omega}_{\bullet{j}},\omega_{jj})^{T}$ in Step 5 of the forward mapping, whereas $\tilde{\boldsymbol{\theta}}_{j}$ is a complicated function of ${\boldsymbol{\theta}}_{j},\ldots,{\boldsymbol{\theta}}_{p}$, as specified in Equation (12). The notational justification for the reverse mapping is similar, as is the justification for a lack of explicit functional form for the functions $\phi(\cdot)$ and $\gamma(\cdot)$ in Equations (12)–(13).
@@
A critical point in Equation (10) is that the partial regression here is of $\mathbf{y}_{p-1}$ on $\mathbf{y}_{1:p-2}$ and not of $\mathbf{y}_{p-1}$ on $\mathbf{y}_{-(p-1)}:=\mathbf{y}_{1:(p-2)\cup p}$. Therefore, the starting point of Equation (10) is the Schur-complement adjusted precision matrix $\boldsymbol{\Omega}_{(p-1)\times(p-1)}$ of Equation (9), and not $\boldsymbol{\Theta}_{(p-1)\times(p-1)}$, as in partial regression based neighborhood selection approaches (Meinshausen and Bühlmann,, 2006), which regresses $\mathbf{y}_{j}$ on $\mathbf{y}_{-j}$ and hence, targets the pseudolikelihood and not the true likelihood. In contrast, in our approach, $\mathbf{y}_{p}$ has been adjusted out when we consider the precision matrix of $\mathbf{y}_{1:p-1}$, as in iterative proportional scaling (Speed and Kiiveri,, 1986). Hence, we target the true likelihood, since by Equation (11), $f(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{1:j-1},\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{1:j-1},\boldsymbol{\theta}_{j},\ldots,\boldsymbol{\theta}_{p})$. However, $f(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})\neq f_{\mathrm{pseudo}}(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{-j},\boldsymbol{\Theta}_{p\times p})$. In a partial regression based approach, this difference is critical for targeting the true posterior vs. a pseudo-posterior. The distinction remains present if one replaces the pseudolikelihood with any other composite likelihood.
@@
Remark 4 (Differences with Cholesky factor based partial regression).
@@
Pourahmadi, (1999) and Daniels and Pourahmadi, (2002) consider the model ${\bf y}\sim\mathcal{N}(0,\boldsymbol{\Sigma})$ and show that it can be written as a sequence of partial regressions: ${\bf y}_{j}=\sum_{k=1}^{j-1}\phi_{kj}{\bf y}_{k}+\varepsilon_{j};\;\varepsilon_{j}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(0,\sigma_{j}^{2}).$ If $\mathbf{T}$ is a unit lower triangular matrix with elements $\phi_{kj}$ and $\mathbf{D}=\mathrm{diag}(\sigma_{j}^{2})$, then Pourahmadi, (1999) establishes that: $\mathbf{T}\boldsymbol{\Sigma}\mathbf{T}^{T}=\mathbf{D}.$ Consequently, Pourahmadi, (1999) proceeds to model $(\phi_{kj},\sigma_{j}^{2})\in\mathbb{R}\times\mathbb{R}_{+}$, with no additional restrictions needed to ensure positive definiteness of $\boldsymbol{\Sigma}$. While this approach shares the similarity with us that the regression is indeed of $\mathbf{y}_{j}$ on $\mathbf{y}_{1:j-1}$, our regression coefficients directly use the elements of $\boldsymbol{\Theta}=\boldsymbol{\Sigma}^{-1}$ and not its Cholesky factor. Thus, we must explicitly enforce positive definiteness of $\boldsymbol{\Theta}$. Apart from specific cases, the zero patterns of $\mathbf{T}$ and $\boldsymbol{\Sigma}$ (or their inverses) do not correspond to each other, and hence, a Cholesky-based approach is not useful for directly modeling a sparse $\boldsymbol{\Theta}$, only a sparse Cholesky factor. Another major difference is that since the Cholesky factors imply a directed acyclic graph (DAG) factorization of the likelihood, inference depends on the ordering of the $p$ variables. In contrast, the proposed method is invariant to such ordering, for the same reason that iterative proportional scaling is invariant to ordering.

### v2

@@ Abstract
We consider the problem of fully Bayesian posterior estimation and uncertainty quantification in undirected Gaussian graphical models via Markov chain Monte Carlo (MCMC) under recently-developed element-wise graphical priors, such as the graphical horseshoe. Unlike the conjugate Wishart family, these priors are non-conjugate; but have the advantage that they naturally allow one to encode a prior belief of sparsity in the off-diagonal elements of the precision matrix, without imposing a structure on the entire matrix. Unfortunately, for a graph with $p$ nodes and with $n$ samples, the state-of-the-art MCMC approaches for the element-wise priors achieve a per iteration complexity of $\mathcal{O}(p^{4}),$ which is prohibitive when $p\gg n$. In this regime, we develop a suitably reparameterized MCMC with per iteration complexity of $\mathcal{O}(p^{3})$, providing a one order of magnitude improvement, and consequently bringing the per iteration computational cost at par with the conjugate Wishart family, which is also $\mathcal{O}(p^{3})$ due to a use of the classical Bartlett decomposition, but this decomposition does not apply outside the Wishart family. Importantly, the proposed benefit is obtained solely due to our reparameterization in an MCMC scheme targeting the true posterior, that reverses the recently developed telescoping block decomposition of Bhadra et al., (2024), in a suitable sense. There is no variational or any other approximate Bayesian computation scheme considered in this paper that compromises targeting the true posterior. Simulations and the analysis of a breast cancer data set confirm both the correctness and better algorithmic scaling of the proposed reverse telescoping sampler.
@@ 1 Introduction
We denote by $p$ the number of variables at hand and by $n$ the sample size. Our specific focus in this article is on Bayesian estimation of $\boldsymbol{\Theta}=\{\theta_{jk}\}$ in the regime $p\gg n$. To begin, we note that this setting is computationally challenging even for routine priors, such as the conjugate Wishart family. Assume one observes $n$ independent and identically distributed (i.i.d.) samples $\mathbf{y}_{i\bullet}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Theta}^{-1})\in\mathbb{R}^{p}$ for $i=1,\ldots,n$ and the prior on $\boldsymbol{\Theta}$ is $\mathcal{W}_{p}(b,\boldsymbol{\Phi})$ for some b>(p1)b>(p-1) and $\boldsymbol{\Phi}\in\mathcal{M}_{p}^{+}$, where $\mathcal{M}_{p}^{+}$ denotes the space of $p\times p$ positive definite matrices, and $\mathcal{W}_{p}(b,\boldsymbol{\Phi})$ denotes a Wishart random variable with shape parameter $b$ and scale matrix $\boldsymbol{\Phi}$ (Gupta and Nagar,, 2000, Chapter 3). Let $\mathbf{y}=(\mathbf{y}_{1\bullet},\ldots,\mathbf{y}_{n\bullet})^{T}\in\mathbb{R}^{n\times p}$. Then, by routine calculations, the posterior of $\boldsymbol{\Theta}$ is also Wishart, given as: $\boldsymbol{\Theta}\mid\mathbf{y}\sim\mathcal{W}_{p}(b+n,\;\boldsymbol{\Phi}+\mathbf{y}^{T}\mathbf{y})$. In this setting, the standard algorithm for posterior sampling is Bartlett’s decomposition of Wishart (Bartlett,, 1934), which requires a Cholesky decomposition of a $p\times p$ matrix, followed by multiplication of two $p\times p$ matrices to generate each draw, and thus has a per sample computational cost of $\mathcal{O}(p^{3})$. Unfortunately, the situation is worse for non-conjugate priors. A crucial bottleneck in the posterior sampling of $(\boldsymbol{\Theta}\mid\mathbf{y})$ under any prior is that one must ensure the posterior samples also belong to $\mathcal{M}_{p}^{+}$. This restriction is automatically satisfied in the conjugate Wishart family, which puts mass only on the space of positive definite matrices, but must be enforced explicitly under all other non-conjugate priors. Therefore, unlike in regression problems with real-valued coefficients, it is far more challenging to put element-wise priors on $\theta_{jk}$ and then to design a valid posterior sampler respecting the constraint $\boldsymbol{\Theta}\in\mathcal{M}_{p}^{+}$. A naïve accept–reject sampler is not likely to be effective in enforcing this constraint at a large enough $p$, because the space $\mathcal{M}_{p}^{+}$ constitutes a very small subset of $\mathbb{R}^{p\times p}$. A Cholesky-based modeling approach, where one writes $\boldsymbol{\Theta}=\mathbf{L}\mathbf{L}^{T}$, is a possibility for maintaining positive definiteness while allowing for unconstrained modeling of $\mathbf{L}$ (Pourahmadi,, 1999; Daniels and Pourahmadi,, 2002), but again, a prior on $\mathbf{L}$ in general induces a complicated prior structure on $\boldsymbol{\Theta}$, and vice versa, apart from special cases such as the Wishart or G-Wishart (Roverato,, 2002; Uhler et al.,, 2018). Thus, if one desires a sparsity-inducing prior directly on $\theta_{jk}$ to learn a sparse graphical structure, it is unclear what induced prior on $\mathbf{L}$ would achieve this goal, apart from very specific cases.
@@
An ingenious approach that indeed allows one to put element-wise non-conjugate priors on $\theta_{jk}$ and still perform Gibbs sampling while respecting the positive definite constraint, was developed by Wang, (2012) in the context of the Bayesian graphical lasso (BGL) prior. We describe this in detail in Section 2. However, the parameterization of Wang, (2012) is not specific to the BGL, and similar sampling approaches have been used later for other priors with better statistical properties, such as the graphical horseshoe or GHS (Li et al.,, 2019), the graphical horseshoe-like or GHSL (Sagar et al., 2024a, ), and the graphical R2-D2 (Gan et al.,, 2022), among others. In particular, Sagar et al., 2024a studied the posterior concentration properties of the element-wise graphical priors and showed that the GHS and GHSL priors enjoy the optimal rates of posterior contraction established by Banerjee and Ghosal, (2014). The penalized maximum a posteriori (MAP) estimate under these element-wise priors have also been studied. For the BGL prior, the MAP corresponds to the frequentist graphical lasso (Friedman et al.,, 2008), whereas a novel local linear approximation scheme for MAP estimation under generic element-wise priors was developed by Sagar et al., 2024b . While MAP estimation is faster, it does not provide full posterior uncertainty quantification, and we do not discuss this further in the current work. Nevertheless, the decomposition of Wang, (2012) is also fundamental to the MAP approaches (Wang,, 2014; Sagar et al., 2024b, ).
@@
Despite its utility, Wang’s parametrization has a fundamental bottleneck, which is a $\mathcal{O}(p^{4})$ complexity per iteration to generate a sample of $(\boldsymbol{\Theta}\mid\mathbf{y})$, which is one order of magnitude worse than the $\mathcal{O}(p^{3})$ complexity for the conjugate Wishart model, and is prohibitive when $p$ is large. We describe in detail in Section 2 why this complexity is inherent to Wang’s approach, and there is no way to bypass this, even if $n$ is much smaller than $p$. Intuitively, this happens because Wang, (2012)’s estimation procedure only sees the $p\times p$ scatter matrix $\mathbf{S}=\mathbf{y}^{T}\mathbf{y}$, defined as $n$-times the sample covariance matrix, and not the $n\times p$ data matrix $\mathbf{y}$ itself. Thus, even if $n$ is much smaller than $p$, there is no way for Wang, (2012) to take advantage of this.
@@
In this work, we demonstrate that a valid Gibbs sampler targeting the exact posterior is possible with a worst case per iteration complexity of $\mathcal{O}(p^{3})$ when $p\gg n$, providing a one order of magnitude reduction in computation cost, and bringing the cost down at par with the conjugate Wishart family, even for nearly arbitrary non-conjugate element-wise graphical priors. The key to our work is to reverse, in a suitable sense, the telescoping block decomposition of the precision matrix recently developed by Bhadra et al., (2024) for estimating the marginal likelihood in GGMs under element-wise priors. However, the possibility of using this decomposition for sampling was not developed by Bhadra et al., (2024), who end up using a $\mathcal{O}(p^{5})$ algorithm. As we demonstrate in the current work, this decomposition is also particularly amenable to fast sampling under Gaussian scale mixture priors, developed in the context of linear regression problems by Bhattacharya et al., (2016). Using this, we demonstrate that our sampler has a per iteration complexity of $\mathcal{O}(\max{(n^{2}p^{2},p^{3})})$, which is $\mathcal{O}(p^{3})$ in the $p\gg n$ regime, i.e., if $n=\mathdutchcal{o}(p^{1/2})$. We take particular care to distinguish our methodology with existing approaches that also use a regression setting, such as the Cholesky factor-based methods (e.g. Pourahmadi,, 1999) or pseudolikelihood methods (e.g. Meinshausen and Bühlmann,, 2006). We perform extensive simulations under the GHS and GHSL priors to verify both the correctness of our algorithm, and its improved scaling over $p$.
@@ 2.3 Model Posterior and the Cyclical Gibbs Sampler of Wang, (2012)
    For the GHS prior of Li et al., (2019), $\lambda_{jk}\sim C^{+}(0,1)$, a priori, where $C^{+}$ denotes the half-Cauchy density. However, introducing one further level of augmentation, one can show: $\lambda_{jk}\sim C^{+}(0,1)$ is equivalent to $\lambda_{jk}^{2}\mid a\sim\mathrm{IG}(1/2,1/a)$ and $a\sim\mathrm{IG}(1/2,1)$, where IG denotes an inverse gamma random variable in shape–scale parameterization. This leads to Gibbs sampling from the posterior as in Makalic and Schmidt, (2016), with $(\lambda^{2}_{jk}\mid\mathrm{rest})$ and $(a\mid\mathrm{rest})$ both following inverse gamma distributions.
@@
Similar strategies are available for posterior sampling under the GHSL prior (Sagar et al., 2024a, ), the graphical R2-D2 prior (Gan et al.,, 2022) and several others. However, it is important to note that the positive definite constraint that applies to the sampling of $\boldsymbol{\Theta}$ does not apply to the sampling of these hyperparameters in a Gibbs step, which is conditional on the current $\boldsymbol{\Theta}$. Finally, for the global $\tau$, a standard half-Cauchy prior is typical (Li et al.,, 2019).
@@ 3.1 The Reverse Telescoping Block Decomposition
Remark 2 (The notations $\boldsymbol{\Theta}$ versus $\boldsymbol{\Omega}$).
We use the notation $\boldsymbol{\Theta}$ and $\boldsymbol{\tilde{\Theta}}$ to denote the matrices, including all their submatrices, that remain unchanged throughout the telescoping and reverse telescoping adjustments and all sampling steps; whereas the notations $\boldsymbol{\Omega}$ and $\boldsymbol{\tilde{\Omega}}$ denote placeholder matrices whose upper-left $(j\times j)$ blocks are successively overwritten as the algorithms proceed over $j=(p-1),\ldots,1$. We clarify this distinction using keywords “Set” versus “Update” in our algorithm description. This allows us to write $\tilde{\boldsymbol{\theta}}_{j}=(\boldsymbol{\omega}_{\bullet{j}},\omega_{jj})^{T}$ in Step 5 of the forward mapping, whereas $\tilde{\boldsymbol{\theta}}_{j}$ is a complicated function of ${\boldsymbol{\theta}}_{j},\ldots,{\boldsymbol{\theta}}_{p}$, as specified in Equation (12). The notational justification for the reverse mapping is similar, as is the justification for a lack of explicit functional form for the functions $\phi(\cdot)$ and $\gamma(\cdot)$ in Equations (12)–(13).
@@
A critical point in Equation (10) is that the partial regression here is of $\mathbf{y}_{p-1}$ on $\mathbf{y}_{1:p-2}$ and not of $\mathbf{y}_{p-1}$ on $\mathbf{y}_{-(p-1)}:=\mathbf{y}_{1:(p-2)\cup p}$. Therefore, the starting point of Equation (10) is the Schur-complement adjusted precision matrix $\boldsymbol{\Omega}_{(p-1)\times(p-1)}$ of Equation (9), and not $\boldsymbol{\Theta}_{(p-1)\times(p-1)}$, as in partial regression based neighborhood selection approaches (Meinshausen and Bühlmann,, 2006), which regresses $\mathbf{y}_{j}$ on $\mathbf{y}_{-j}$ and hence, targets the pseudolikelihood and not the true likelihood. In contrast, in our approach, $\mathbf{y}_{p}$ has been adjusted out when we consider the precision matrix of $\mathbf{y}_{1:p-1}$, as in iterative proportional scaling (Speed and Kiiveri,, 1986). Hence, we target the true likelihood, since by Equation (11), $f(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{1:j-1},\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{1:j-1},\boldsymbol{\theta}_{j},\ldots,\boldsymbol{\theta}_{p})$. However, $f(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})\neq f_{\mathrm{pseudo}}(\mathbf{y}\mid\boldsymbol{\Theta}_{p\times p})=\prod_{j=1}^{p}f(\mathbf{y}_{j}\mid\mathbf{y}_{-j},\boldsymbol{\Theta}_{p\times p})$. In a partial regression based approach, this difference is critical for targeting the true posterior versus a pseudo-posterior. The distinction remains present if one replaces the pseudolikelihood with any other composite likelihood.
@@
Remark 4 (Differences with Cholesky factor-based partial regression).
@@
Pourahmadi, (1999) and Daniels and Pourahmadi, (2002) consider the model ${\bf y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}\otimes\boldsymbol{\Sigma})$ and show that it can be written as a sequence of partial regressions: ${\bf y}_{j}=\sum_{k=1}^{j-1}\phi_{kj}{\bf y}_{k}+\varepsilon_{j};\;\varepsilon_{j}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(0,\sigma_{j}^{2}\mathbf{I}_{n}).$ If $\mathbf{T}$ is a unit lower triangular matrix with elements $\phi_{kj}$ and $\mathbf{D}=\mathrm{diag}(\sigma_{j}^{2})$, then Pourahmadi, (1999) establishes that: $\mathbf{T}\boldsymbol{\Sigma}\mathbf{T}^{T}=\mathbf{D}.$ Consequently, Pourahmadi, (1999) proceeds to model $(\phi_{kj},\sigma_{j}^{2})\in\mathbb{R}\times\mathbb{R}_{+}$, with no additional restrictions needed to ensure positive definiteness of $\boldsymbol{\Sigma}$. While this approach shares the similarity with us that the regression is indeed of $\mathbf{y}_{j}$ on $\mathbf{y}_{1:j-1}$, our regression coefficients directly use the elements of $\boldsymbol{\Theta}=\boldsymbol{\Sigma}^{-1}$ and not its Cholesky factor. Thus, we must explicitly enforce positive definiteness of $\boldsymbol{\Theta}$. Apart from specific cases, the zero patterns of $\mathbf{T}$ and $\boldsymbol{\Sigma}$ (or their inverses) do not correspond to each other, and hence, a Cholesky factor-based approach is not useful for directly modeling a sparse $\boldsymbol{\Theta}$, only a sparse Cholesky factor. Another major difference is that since the Cholesky factors imply a directed acyclic graph (DAG) factorization of the likelihood, inference depends on the ordering of the $p$ variables. In contrast, the proposed method is invariant to such ordering, for the same reason that iterative proportional scaling is invariant to ordering.